OpenStudy (anonymous):

How do you simplify z^-20 multiplied by z^3

OpenStudy (anonymous):

Hint: 

ganeshie8 (ganeshie8):

simply add the exponents \[\large z^{-20}\cdot z^3 = z^{-20+3}=?\]

ganeshie8 (ganeshie8):

then use the above hint

OpenStudy (anonymous):

Thank you!